(x+2)^2+28=0 solve using square root procedure
Answer by stanbon(75887)

(x+2)^2+28=0 solve using square root procedure
-----
(x+2)^2 = -28
---
x+2 = +/-2isqrt(7)
---
x = -2 + 2isqrt(7) or x = -2 - 2isqrt(7)
===========================================
Cheers,
Stan H.
===================